# thanks: These authors contributed equallythanks: These authors contributed equallythanks: These authors contributed equallythanks: These authors contributed equally

### v1

@@ IX Conclusions
In this connection, we have incorporated in the EPW code the computation of off-diagonal mode-mode couplings in the phonon self-energy, which has enabled us to uncover the role of anti-crossing phenomenon in generating Kohn anomalies. An in-depth study of the Giant Kohn anomaly and chiral phonons in 1H-NbSe2 transition-metal-dichalcogenide is presented as an illustrative application of our modified EPW code. We provide compelling evidence that the phonon that softens in NbSe2 is a longitudinal optical phonon which softens by anti-crossing several intervening phonon bands. We also show that the CDW vector in NbSe2 results from the convolution of the susceptibility and electron-phonon coupling, and that the softened phonons are circularly polarized. These chiral phonons may be relevant for generating time crystals, reducing decoherence in solid-state defect-spin qubits, among other potential applications. Our study will impact the modeling of charge density waves and other phenomena involving electron-phonon couplings by enabling accurate, material-specific computation of phonon self-energies in wide classes of materials.
Appendix A Deconvolving the self-energy
To better understand the origin of the two peaks at $q_{1}$ and $q_{2}$, found in $\Pi$, we plot the $k$-resolved contributions to $\Pi(\Omega_{q})$ in Fig.A.1. Panel (a) shows electronic dispersion over the full Brillouin zone and the Fermi surface to determine corresponding vectors for the peaks. In panel (c), we plot $|g_{k,{k+q}}|^{2}$, in panel (b) $\chi_{q,k}$, the integrand of Eq. 7, and in panel (b) we superpose the two plots. We have confirmed that the integral of this superposition over the Brillouin zone gives the calculated values of $\Pi$ at those $q$-values. We note that the $\chi$ maps measure nesting – the overlap of $\epsilon_{k}$ and $\epsilon_{k+q}$, while the $G$ maps tend to enhance features in approximately the same $k$-regions of the susceptibility maps, where the nesting is significantly better for $q_{2}$. In Fig. A.2 we compare the clean and dirty limit self energies.
Figure A.1: (color online) Deconvolution of self-energy, (a)Left: Energy dispersion over the full Brillouin zone, with energy increasing from dark blue to dark red, Right: Fermi surface, (b)-(d) are respectively $G(k)\cdot\chi(k)$, $G(k)=|g_{\bf k,\bf k+\bf q}|^{2}$ and $\chi(k)$. Left figures are for peak between $\Gamma M$ (peak $q_{1}$) and Right ones are for the peak between $MK$ (peak $q_{2}$)
Figure A.2: (color online) Phonon spectra and self energies at three effective temperatures, from left to right $\sigma$ = 0.0057, 0.009, 0.02 Ry. Row (a) phonon modes, (b) self-energy in clean limit, and (c) self-energy in dirty limit. All frames use same color codes to identify the different phonon modes.
Appendix B Density functional perturbation theory (DFPT) and self-energies
The dressed phonon frequencies $\Omega_{q}$ are calculated in QE directly from density functional perturbation theory (DFPT) [Baroni]. To better understand the ‘bare’ phonon frequencies we have found, we briefly recapitulate the derivation of the QE formalism[QuantEspRenorm].
The physical basis of these calculations is the Born-Oppenheimer approximation, where the nuclear degrees of freedom can be separated from electronic degrees of freedom. Within the harmonic approximation, the phonon spectra and eigenmodes can be calculated using the Hessian of the BO-energy surface scaled by the geometric mean of the masses of the nuclei
$\left(\frac{1}{\sqrt{M_{I}M_{J}}}\frac{\partial^{2}E(R)}{\partial R_{I}\partial R_{J}}-\omega^{2}\mathbf{I}_{3N\times 3N}\right)\mathbf{\hat{u}}=0,$ (13)
where the potential energy surface $E(R)$ is calculated using the BO Hamiltonian for electrons, and $\mathbf{I}_{3N\times 3N}$ is the identity matrix. The non-trivial solutions give the eigenfrequencies $\omega_{\mu}$ and eigenvectors $\mathbf{\hat{u}}_{\mu}$ for phonons. In Baroni et al. [Baroni], the Hellman-Feynman theorem and BO-Hamiltonian are applied, which leads to a useful form of of Hessian:
[…]
where the $\mathbf{F}_{I}$ are the interatomic forces and $V_{R}(r)$ is the electron-nucleus interaction.
Equation 14 is the chief result of this Appendix. While not immediately obvious, the first term
[…]
contains a static coupling between electrons and phonons, since using linear response theory (see, e.g., Baroni Eq. (69) [Baroni]) the first order variation of the electron density in terms of Kohn-Sham eigenenergies and eigenfunctions indicates the appearance of a non-retarded (static) response function
[…]
When Fourier transformed and transformed into the basis of phonon modes, this corresponds to the static limit as Equation 9. Thus, the first term of Eq. 14 corresponds to our $\Pi^{\prime}$, so the remaining terms ${\bf\Pi}^{B}$ constitute $\omega_{q}^{2}$. We note that one term is the potential of the static ions, while the other is the interaction of the ions with the average electron density $\left\langle n_{R}\right\rangle$, which has a temperature dependence via the Fermi function. Moreover, since the standard DFpT calculation in QE includes the static part of the charge density variation, this term includes the main difference between $\Omega_{q}^{2}$ and the “bare” phonons $\omega_{q}^{2}$, using the definition of Ref. [Giustino1].
To further relate our results to current literature, we note the following: $\Pi^{\prime}_{\lambda,\lambda^{\prime}}(\boldsymbol{q})$ is the adiabatic self-energy as defined by Giustino [Giustino1]. If this is subtracted from the frequency dependent self-energy of Eq. 9 one obtains a non-adiabatic self-energy term with a retarded response of the electronic degrees of freedom to displacements of nuclei. In EPW the DFpT phonons are the bare phonons which are dressed by the non-adiabatic self energy to produce the dressed or dynamical phonons. More specifically, the correction is assumed to be small enough that only the diagonal self energy is important, so the non-adiabatic self energy is defined as $\Pi^{NA^{\prime}}(q,\Omega_{q})=\Pi^{\prime}_{\lambda,\lambda}({\boldsymbol{q}},\omega=\Omega_{\lambda})-\Pi^{\prime}_{\lambda,\lambda}(\boldsymbol{q})$ (Eq. 145 in Ref. Giustino1). In fact, adiabatic and non-adiabatic self-energy terms can be equated to static and dynamical terms of self-energy, as explicitly discussed in Ref. [ultrasoft], where they make a distinction between static and dynamical phonons ((bare and dressed phonons in the terminology of Ref. [Giustino1]).
However, we find that this correction is small, and the main softening is contained in the DFPT phonons, necessitating our rewriting of the dynamical matrix to find the bare BO phonons – i.e., without the e-ph coupling self energy.
Finally, we note that in Eq. 14 the unknown $\omega$ arises only in the $\omega^{2}$ term, inconsistent with Eq. 17, but consistent with Eq. 10. This supports our idea that in the term $\omega\Pi$, the prefactor $\omega$ exactly cancels the normalization-derived $1/\omega$ in $\Pi$, leaving $\Pi^{\prime}$ the essential quantity to calculate.
Appendix C Inverse dynamical matrix
In trying to find the bare phonon frequencies in terms of the dressed phonons, there are two possible choices. The simplest procedure is to modify Eq. 1 by replacing $\omega_{\boldsymbol{q}}$ by $\Omega_{\boldsymbol{q}}$ in the self-energy $\Pi$,
$\det\Bigl{\{}({\boldsymbol{\Omega}}_{\boldsymbol{q}}^{2}-{\omega^{2}})\operatorname{\boldsymbol{I}}-2\,{\omega}\,{\boldsymbol{\Pi}}(\boldsymbol{q},{\boldsymbol{\Omega}}_{\boldsymbol{q}})\Bigr{\}}=0.$ (17)
Since the unknown $\omega$ occurs explicitly in the dynamical matrix, the bare frequencies are not the solutions of a determinant equation, but must be found self-consistently. However, if the off-diagonal elements of $\Pi$ are neglected, Eq. 17 becomes a quadratic equation. The inverse solution to Eq. 17 is
$\omega_{q}=-\Pi(\Omega_{q})+\sqrt{\Omega_{q}^{2}+\Pi^{2}(\Omega_{q})}\rightarrow-2\Pi(\Omega_{q}),$ (18)
where the limiting behavior holds near the soft mode, $\Omega_{q}=0$ (recall that $\Pi$ is negative). Thus, Eq. 17 leads to an unphysical divergence of $\omega_{q}$ every time there is a soft mode, Fig. C.1. We believe that this is a shortcoming of the Hedin approach, in that for an equation like Eq. 1 it is not clear which $\omega_{q}$-terms remain, and which are replaced by $\Omega_{q}$. Here, we propose an alternative solution. Noting that in the term $\Pi^{\prime}(\omega_{q})=\omega_{q}\Pi(\omega_{q})$, the prefactor $\omega_{q}$ exactly cancels the normalization factor $1/\omega_{q}$ in $\Pi$, we propose that this should be an exact result. Equivalently, in the Hedin-like expression $\Pi^{\prime}(\omega_{q})$ should be replaced by $\Pi^{\prime}(\Omega_{q})$. When this is done, Eq. 17 becomes
$\det\Bigl{\{}(\Omega^{2}-{\boldsymbol{\omega}_{\boldsymbol{q}}^{2}})\operatorname{\boldsymbol{I}}-2\Omega{\boldsymbol{\Pi}}(\boldsymbol{q},\Omega)\Bigr{\}}=0,$ (19)
which can be reorganized to Eq. 10. Note that Eq. 10 can be thought of as the equation for the zeroes of an anti-Green’s function which derives the bare frequencies from the dressed. Neglecting off-diagonal self energies, the solution becomes
$\omega_{q}^{2}=\Omega_{q}^{2}-2\Omega_{q}\Pi(\Omega_{q})\rightarrow-2\Pi^{\prime}(\Omega_{q}).$ (20)
In Fig. 4 we show the bare phonon frequencies calculated from Eq. 20 as dashed lines. These results should be contrasted with Fig. C.1.
Figure C.1: (color online) Single mode calculations for the three longitudinal modes, modes 3, 6, and 8. Solution of Eq. 18 for $\sigma$ =0.057 Ry, comparing the dressed frequencies $\Omega_{q}$ (dashed curves) with the bare frequencies $\omega_{q}$ (solid curves).
Appendix D DFT Parameters and Potentials
@@
Density functional theory (DFT) calculations were performed using the Quantum ESPRESSO package. The structural optimization was carried out for mono-layer 1H-NbSe2 by including ${\AA}$ and setting the energy and force convergence thresholds to $10^{-6}$ and $10^{-5}$ Ry, respectively. A plane-wave basis set was employed, with a wavefunction cutoff of 100 Ry and a charge density cutoff of 800 Ry. For the electronic structure, a convergence threshold of $10^{-12}$ Ry was used, incorporating local Thomas-Fermi screening. Fermi-Dirac smearing was applied to model temperature effects, and the smearing value was varied accordingly. Ultrasoft pseudopotentials were used for Nb, while norm-conserving pseudopotentials were selected for Se, following the methodology outlined in Ref. anharm. Self-consistent field (SCF) calculations were performed using a $36\times 36\times 1$ Monkhorst-Pack $K$-grid. Phonon spectra were computed with a precision of $10^{-16}$ Ry on a $12\times 12\times 1$ mesh. The electron-phonon coupling (EPC) matrix was evaluated using the EPW code along a $Q$-path sampled with 500 points, utilizing a dense $500\times 500\times 1$ $k$-point mesh.
Appendix E Kekule-like Graphene
Figure E.1: (color online) K-point phonon displacements in graphene. Top and bottom rows show the same three phases of a single cycle of the highest K-point phonon in graphene.
Using Phonon Visualizer, we find similar Kekule-like phonon displacement patterns in benzene, Fig. E.1. The top row focuses on a breathing-in mode for a single hexagon [yellow squares], as it cycles among three possible centers for the hexagon. In the bottom row, we focus instead on a single hexagon [purple squares] as it evolves through the same cycle. We observe the same Kekule-like pattern as seen in NbSe2.
@@ Acknowledgements
The work at Northeastern University was supported by the National Science Foundation through the Expand-QISE award NSF-OMA-2329067 and benefited from the resources of Northeastern University’s Advanced Scientific Computation Center, the Discovery Cluster, the Massachusetts Technology Collaborative award MTC-22032, and the Quantum Materials and Sensing Institute. S.C. and S.E. acknowledge the support provided by the National Science Foundation through the ExpandQISE award No. 2329087. S.C. and S.M. are supported by the U.S. Department of Energy (DOE), Office of Science, and Basic Energy Sciences under Award No. DE-SC0022216. The calculations performed by T.H. and C.S. in this paper were performed at TUBITAK ULAKBIM, High Performance and Grid Computing Center (TRUBA resources). The research at Howard University used the resources of Accelerate ACCESS PHYS220127 and PHYS2100073. We thank Adrian Feiguin for stimulating discussions. T.H. is thankful to Cagliyan Kurdak and the Department of Physics of the University of Michigan for partial support.

### v2

@@ 2 Conclusions
We have developed a technique to uncover the quantum strangeness associated with anti-crossing phenomena and their important role in the softening of optical phonons. An in-depth study of the giant Kohn anomaly and circular phonons in 1H-NbSe2 transition-metal-dichalcogenide is presented as an illustrative application of our technique. We provide compelling evidence that the soft phonon in NbSe2 is a longitudinal optical phonon which softens by anti-crossing several intervening phonon bands. We also show that the CDW vector in NbSe2 results from the convolution of the susceptibility and electron-phonon coupling, and that the softened phonons are circularly polarized. Such phonons may be relevant for generating time crystals, reducing decoherence in solid-state defect-spin qubits, among other potential applications. Our calculations required an extension of the EPW code to compute off-diagonal mode-mode couplings in the phonon self-energy. Our study will impact the modeling of charge density waves and other phenomena involving electron-phonon couplings by enabling accurate, material-specific computation of phonon self-energies in wide classes of materials.
3 Methods
@@
We performed first-principles calculations based on density functional theory (DFT) using the Quantum ESPRESSO package 29, 28. The monolayer 1H-NbSe2 structure was optimized with a vacuum spacing of 10 Å. We used a plane-wave basis set with a wavefunction cutoff of 100 Ry and a charge density cutoff of 800 Ry. The energy and force convergence thresholds were set to $10^{-6}$ and $10^{-5}$ a.u., respectively. We employed ultrasoft pseudopotentials for Nb and norm-conserving pseudopotentials for Se, following the methodology from the literature. 9. The self-consistent field (SCF) calculations were performed on a $36\times 36\times 1$ Monkhorst-Pack $k$-point grid.
Phonon spectra and electron-phonon coupling matrix elements were calculated using density functional perturbation theory (DFPT) 6. The phonon frequencies were computed on a $12\times 12\times 1$ $q$-point mesh with a precision of $10^{-16}$ Ry. The electron-phonon coupling (EPC) matrix was evaluated using the EPW code 31 along a $Q$-path sampled with 500 points, utilizing a dense $500\times 500\times 1$ $k$-point mesh. To account for temperature effects, we used Fermi-Dirac smearing with a broadening parameter $\sigma$, related to an effective temperature $T_{\sigma}=\sigma/k_{B}$.
To investigate the role of phonon mode coupling, we extended the EPW code to compute the off-diagonal elements of the phonon self-energy $\Pi_{\lambda\lambda^{\prime}}$. The adiabatic self-energy $\Pi^{\prime}_{\lambda\lambda^{\prime}}$ was calculated as defined in the cited paper. 31. The dressed phonon frequencies $\Omega_{q}$ were obtained from DFPT, and the bare phonon frequencies $\omega_{q}$ were extracted by inverting the dynamical matrix, as discussed in the Supporting Information (see LABEL:SM3 and LABEL:SM4).
@@ Acknowledgements
The work at Northeastern University was supported by the National Science Foundation through the Expand-QISE award NSF-OMA-2329067 and benefited from the resources of Northeastern University’s Advanced Scientific Computation Center, the Discovery Cluster, the Massachusetts Technology Collaborative award MTC-22032, and the Quantum Materials and Sensing Institute. S.C. and S.E. acknowledge the support provided by the National Science Foundation through the ExpandQISE award No. 2329087. S.M. was supported by the U.S. Department of Energy (DOE), Office of Science, and Basic Energy Sciences under Award No. DE-SC0022216. The calculations performed by T.H. and C.S. in this paper were performed at TUBITAK ULAKBIM, High Performance and Grid Computing Center (TRUBA resources). The research at Howard University used the resources of Accelerate ACCESS PHYS220127 and PHYS2100073. We thank Adrian Feiguin for stimulating discussions. T.H. is thankful to Cagliyan Kurdak and the Department of Physics of the University of Michigan for partial support.
4 Associated Content
A preprint version of this work is available as Susy Exists, Sougata Mardanya, Robert Markiewicz, Tugrul Hakioglu, Jouko Nieminen, Ville J. Härkönen, Cem Sanga, Arun Bansil, and Sugata Chowdhury, Giant Kohn anomaly and chiral phonons in the charge density wave phase of 1H-NbSe2, 2025, arXiv:2501.04821, arXiv, https://doi.org/10.48550/arXiv.2501.04821 (accessed 7 January, 2026).
Supporting Information Available: Toy model of Kohn ladders; k-resolved deconvolution of the phonon self-energy; DFPT/self-energy formalism; inverse dynamical matrix approach for extracting bare phonon frequencies; DFT/DFPT/EPC computational parameters and pseudopotentials; Kekulé-type $K$-point phonon displacement patterns in graphene.